how do i find the completion of squares for the equation y2-5y+6=0
Found 2 solutions by ewatrrr, stanbon:
Answer by ewatrrr(24785)


Hi,
y^2-5y+6=0
Factoring
(y-3)(y-2) = 0 Note inner product(-3y) and outer product (-2y) have a sum of -5y
(y-3)=0
y = 3
(y-2) = 0
y = 2

Answer by stanbon(75887)

how do i find the completion of squares for the equation
y^2-5y+6=0
----
y^2-5y+? = -6+?
-----
Complete the square:
y^2-5y+(5/2)^2 = -6+(5/2)^2
---
Factor the left side:
(y-(5/2))^2 = (-24/4)+(25/4)
------
(y-(5/2))^2 = = 1/4
---
y-(5/2) = 1/2 or y-(5/2) = -1/2
----
y = 3 or y = 2
===============================
Cheers,
Stan H.